OpenStudy (anonymous):

solve each equation system by the addition method 3x+4y=7 2x+y=3

OpenStudy (anonymous):

3x+4y=7 2x+y=3<- multiply by 4

OpenStudy (anonymous):

that is 8x + 4y = 12

OpenStudy (anonymous):

8x + 4y = 12 3x + 4y = 7

OpenStudy (anonymous):

add them while adding the sign in the equation below changes

OpenStudy (anonymous):

i.e 3x becames -3s 4y become -4y and 7 becomes -7 This was the equation we added

OpenStudy (anonymous):

so now 9x + 4y = 12 -3x - 4y = -7 agree?

OpenStudy (anonymous):

4y gets canceled

OpenStudy (anonymous):

OMG! OTS 8X AND NOT 9X SORRY MY BAD!

OpenStudy (anonymous):

so now 8x + 4y = 12 -3x - 4y = -7 agree?

OpenStudy (anonymous):

thats 5x = 5

OpenStudy (anonymous):

divide both sides by 5

OpenStudy (anonymous):

that gives x =1

OpenStudy (anonymous):

put x = 1 in 3x + 4y = 7 to get the value of y

OpenStudy (anonymous):

3(1) + 4y = 7 3 + 4y = 7 4y = 4

OpenStudy (anonymous):

divide both sides by 4

OpenStudy (anonymous):

y = 1

OpenStudy (anonymous):

therefore ordered pair is (1,1)

OpenStudy (triciaal):

this is correct but to keep things simple remember the objective is to eliminate one of the variable when you add the equations for this you would multiply by -4 so that the y's cancel less chances of error when trying to switch signs and without explanation

OpenStudy (triciaal):